OpenStudy (anonymous):

A circle has a diameter of twelve millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. thirty-seven point seven millimeters; four hundred fifty-two point two square millimeters eighteen point eight millimeters; one hundred thirteen point zero square millimeters thirty-seven point seven millimeters; one hundred thirteen point zero square millimeters eighteen point eight millimeters; four hundred fifty-two point two square millimeters

terenzreignz (terenzreignz):

All under the assumption that \[\large \pi = 3.14\] Use these formulas... \[\Large Circumference = d \color{red}\times \pi\] \[\Large Area = \frac14 \color{red}\times d\color{red}\times d\color{red}\times\pi\]

OpenStudy (anonymous):

i dont get it

terenzreignz (terenzreignz):

d is the diameter.

terenzreignz (terenzreignz):

So, for example, if our diameter was 5, then the circumference would be... \[\Large Circumference = 5 \color{red}\times \pi\]\[\Large Circumference =5\color{red}\times3.14 \]\[\Large Circumference = 15.7\]

OpenStudy (anonymous):

so 15.7xd

terenzreignz (terenzreignz):

Nope... that was an example.

OpenStudy (anonymous):

oh... okay.....

terenzreignz (terenzreignz):

Here, your diameter is 9 so just take d = 9 and solve.

OpenStudy (anonymous):

okay thanks bye

terenzreignz (terenzreignz):

Wait no it isn't.

terenzreignz (terenzreignz):

Sorry, d = 12 not 9 I'm getting drowsy...